OpenStudy (anonymous):

Given that 1 − 2i is a root of z4 + 11z2 − 10z + 50 = 0 find the other roots.

OpenStudy (joes):

we know that 1+2i is also a root since the coefficients of the polynomial are real. Therefore z^4 + 11z^2 - 10^z +50 = (z^2 - 2z + 5)(Az^2 + Bz +C) By inspection (comparing coefficients) we find A = 1, B = 2, C = 10 therefore solving the equation: (z^2 - 2z + 5)(z^2 - 2z + 10) = 0 is the same as solving z^4 + 11z^2 - 10^z +50 = 0. Im sure you will be able to get the final answer from here